OpenStudy (anonymous):

graph the exponential model y=2(0.5)^x .................which is on that graph up there (2,5) (1,-1) (-4,1) or (0,2)

OpenStudy (anonymous):

You should plug in points: (0.5)^5*2 = 1/16 (0.5)^-1*2 = 4 (0.5)^1*2 = 1 (0.5)^0*2 = 2 D.